OpenStudy (anonymous):

What is the area bounded by the curves y = x^2 and y=the square root of x

OpenStudy (amistre64):

do you know the interval where they intesect?

OpenStudy (amistre64):

\[\int_{a}^{b}f-g~dx\]

OpenStudy (whpalmer4):

Solve \[x^2=\sqrt{x}\] to find the endpoints of the integral.

OpenStudy (anonymous):

I don't understand how to do this, I need help working this out so I can use this as an example for the rest of my work

OpenStudy (amistre64):

this is the help ... define the interval, and integrate the difference of the functions inbetween

OpenStudy (amistre64):

if you get a negative value at the end, drop the sign

OpenStudy (amistre64):

\[\left|\int_{a}^{b}x^2-x^{1/2}~dx\right|\]

OpenStudy (whpalmer4):

 You find the two points where the curves cross to determine the endpoints of the integral. The integral adds up all the infinitesimal rectangles from the function with the greater y value down to the function with the smaller y-value between the endpoints.